A coil with an inductance of 1.9 H and a resistance of 14 Ω is suddenly connected to an ideal battery with ε = 77 V. At 0.14 s after the connection is made, what is the rate at which (a) energy is being stored in the magnetic field, (b) thermal energy is appearing in the resistance, and (c) energy is being delivered by the battery?
Answer #1

L = 1.9H

R = 14 Ω

ε = 77 V

Current I after time t = I = I _{0} ( 1 - e^ -t/ \tau ) where I _{0} = ε /R and \tau =L/R

a) The rate at which energy is stored in the inductor in the form of magnetic potential energy is given as:

P = Vi = LI dI/dt

I after 0.14 s

I = ε /R (1- e^-t/\tau ) = 77/14 (1- e ^ { - 0.14 / (1.9/14) } ) = 3.53977 A

dI/dt = (ε\tau /R) e^-t/\tau = (77/14 * 7.36) * 0.356 = 14.41

P = LI dI/dt = 1.9H * 3.53977 A * 14.41 = 96.89 J/s

Answer is 96.89 J/s

b) Thermal Energy is I^2 R = (3.53977)^2 * 14 = 12.524 * 14 = 175.3 J/s

c) Energy delivered by battery = 96.89 J/s + 175.3 J/s = 272.23 J/s